# GlobalMapNet: An Online Framework for Vectorized Global HD Map Construction

### v1

@@ IV-B GlobalMapBuilder
Map NMS. To further improve the quality of global map construction, we propose a novel post-processing method, namely Map NMS, to remove duplicate predictions of map elements. Similar to NMS in object detection, $\left\{P_{i}^{G^{+}}\right\}$ is first sorted by confidence score, and a map element with higher score eliminates another if their Intersection over Union (IoU) is above the given threshold. Here, buffered IoU [35] is employed to formulate the overlapping between point sequences. With Map NMS, overlap within the same category can be eliminated to produce a clean global map.
@@ V-B Results
The results are shown in Table IV. GlobalMapNet is still much better than StreamMapNet in mGAP (+4.7), confirming its superiority in real scenarios. The enhanced advantage in $\mathrm{GAP}_{ped}$ (+7.0) proves that GlobalMapFusion can benefit from cross-scene information in generating small and static map elements.
@@ V-C Ablation studies
    StreamMapNet is the original baseline model. It utilizes BEV feature propagation and Query propagation, which brings 2.8 mAP increase and 1.3 mGAP increase.
@@
    GlobalMapNet-- replaces Query propagation with GlobalMapFusion. This step brings 1.1 mAP increase and 3.4 mGAP increase.
@@
    GlobalMapNet further utilizes traced region information in map fusion, which brings 0.8 mAP increase and 0.3 mGAP increase.

### v2

@@ IV-B GlobalMapBuilder
Map NMS. To further improve the quality of global map construction, we propose a novel post-processing method, namely Map NMS, to remove duplicate predictions of map elements. Similar to NMS in object detection, $\left\{P_{i}^{G^{+}}\right\}$ is first sorted by confidence score, and a map element with the higher score eliminates another if their Intersection over Union (IoU) is above the given threshold. Buffered IoU [35] is employed to formulate the overlap between point sequences. With Map NMS, overlap within the same category can be eliminated to produce a clean global map.
@@ V-B Results
The results are shown in Table IV. GlobalMapNet is still much better than StreamMapNet in mGAP (+4.7), confirming its superiority in real scenarios. The enhanced advantage in $\mathrm{GAP}_{ped}$ (+7.0) proves that GlobalMapFusion can benefit from cross-scene information in generating small map elements like pedestrian crossings consistently.
@@ V-C Ablation studies
    StreamMapNet is the original baseline model. It utilizes BEV feature propagation and Query propagation, which bring a 2.8 mAP increase and a 1.3 mGAP increase.
@@
    GlobalMapNet-- replaces Query propagation with GlobalMapFusion. This step brings a 1.1 mAP increase and a 3.4 mGAP increase.
@@
    GlobalMapNet further utilizes traced region information in map fusion, which bring a 0.8 mAP increase and a 0.3 mGAP increase.